# On the class of coding optimality of human languages and the origins of Zipf’s law

### v3

@@ 1 Introduction
Table 1: A non-singular coding of numbers from 1 to 6 with an alphabet formed by ${a,b}$.
Unit Code
$1$ $aa$
$2$ $ab$
$3$ $a$
$4$ $b$
$5$ $ba$
$6$ $bb$
Table 2: A coding of numbers from 1 to 6 that is not non-singular.
Unit Code
$1$ $aa$
$2$ $aa$
$3$ $a$
$4$ $b$
$5$ $ba$
$6$ $bb$
In standard information theory, one assumes that $l(i)$ is the length of the code assigned to unit $i$. The code is a string of characters from some alphabet of size $N$. The problem of compression consists of finding the codes whose length minimizes $l\right>$. We use $l\right>_{min}$ to refer to the minimum value of $l\right>$. If no constraint is imposed on the possible codes, the minima of $l\right>$ are singular. If $l(i)=0$ is allowed, $l\right>$ is minimized by assigning the empty string to each unit. Then $l\right>_{min}=0$ and information transmission is impossible with empty strings. If $l(i)>0$ is required, $l\right>$ is minimized by assigning a string of just one character to each unit. Then $l\right>_{min}=1$ and $N$ has to be as large as the number of units for successful decoding, which, if possible, is eventually an overkill. For this reason, information theory is concerned about constraints on the mapping of units into codes, which are known as coding schemes [11]. In the non-singular coding scheme, every unit is assigned a distinct string. Suppose that $N=2$ and the alphabet is $\{a,b\}$. Table 1 shows a coding table that is non-singular while Table 2 shows a coding table that is not non-singular because the string $aa$ is assigned to both unit 1 and unit 2. In the uniquely decodable coding scheme, a particular case of non-singular coding, the segmentation of a string resulting from the concatenation of codes must be unique given the coding table. For instance, $baba$ allows for many interpretations according to Table 1: the segmentation $ba$, $ba$ implies that the units coded are $5,5$; the segmentation $b$, $a$, $b$, $a$ implies that the units coded are $4$,$3$,$4$,$3$, and so on. Therefore, Table 1 does not satisfy the uniquely decodable coding scheme. In contrast, Table 3 does so by using Elias gamma encoding [12]: the sequence $baba$ can only be segmented as $b,aba$ and then the units coded are certainly $1,2$.
@@
Table 3: Elias gamma codes for numbers from 1 to 6 with an alphabet formed by ${a,b}$.
Unit Code
$1$ $b$
$2$ $aba$
$3$ $abb$
$4$ $aabaa$
$5$ $aabab$
$6$ $aabba$
@@ 2 A new class of coding optimality
The following property indicates that, when Zipf’s law holds, the optimal code length under unique decodability is shifted linearly above the optimal code length under non-singular coding. By the same token, $H$ is shifted linearly above $\log_{N}i\right>$. The same applies to the coding efficiencies.
@@
Zipf’s rank-frequency law with exponent $\alpha$ and normalization factor $c$ is equivalent to a linear dependence between optimal code lengths with parameters $\alpha$ and $\beta=-\log_{N}c$ and $\alpha,\beta\geq 0$. First, $l_{min}^{ud}(i)$ and ł$\l_{min}^{ud}(i)$ are linearly dependent as
$l_{min}^{ud}(i)=\alpha\l_{min}^{ud}(i)+\beta.$
@@
(Eqs. 9, 8 and $\sum_{i}p(i)=1$)
Diving by $l\right>$ on the last result, one obtains Eq. 11. We have $\alpha\geq 0$ by definition of rank. To conclude that $\beta\geq 0$, one does not need to know the precise definition of $c$. Notice that $c=p(1)$ (Eq. 7). As $0<p(1)\leq 1$, $\beta>0$ follows. ∎
@@ 5.1 A new interpretation of rank distributions in log-log scale
Our information theoretic framework provides a new insight into plots of frequency versus rank in double logarithmic scale: the appearance of a straight line indicates (linear) proximity between the lengths of two kinds of optimal codes for a word type given its frequency and its rank (Property 1). The two kinds of optimal codes are the non-singular codes and a subclass that are the uniquely decodable codes. The exponent of Zipf’s law is the slope of the linear transformation that links the optimal lengths. In a linguistically diverse sample of languages, $\alpha$ ranges from 0.765 in Thai to 1.442 in Korean [17]. Such an insight does not make any assumption: it just follows from the mathematical shape of Zipf’s law and the optimal lengths of the codes.
A more powerful insight can be obtained by assuming that the system is compressing, namely reducing the length of codes. In that case, the straight line in log-log scale would indicate proximity of the system to optimal coding. If the gap between the length of words and the corresponding optimal length is only linear, then the system must exhibit Zipf’s law (Property 2). Languages (at least Catalan, Spanish and English) exhibit such a linear proximity, which is consistent with a strong form of compression. The failure to find Zipf’s law in certain species but an exponential distribution instead [23, 25, 24] suggests that such a pressure may not be strong enough for these species or that we are not looking at the right units. In humans, part-of-speech tags, word orders, colors, kinship terms, and verbal alternation classes also exhibit an exponential distribution [34, 35, 27, 36].
@@ 5.3 The origins of Zipf’s law
which implies that rank distributions are channeled to show a power-law (Eq. 7) with $\alpha=1$. However, we have reviewed above multiple real examples of systems that exhibit an exponential rank distribution. Another question is if random typing could be the reason why human languages and other systems show Zipf’s law. In the past, random typing has been regarded as a proof that Zipf’s law can be retrieved without involving any optimization [20, 21, 41, 42]. That argument is flawed because random typing is an optimal non-singular coding system. However, one can still retain that random typing is a proof that Zipf’s law is easy to retrieve by virtue of some simple stochastic process. However, this raises the question of why certain species do not exhibit Zipf’s law for word frequencies. The only answer that random typing provides is that $N=1$ yields an exponential distribution while $N>1$ yields a rank distribution and a frequency spectrum whose ability to mimic the actual distribution of word frequencies has been criticized heavily [43, 44].
@@
where $\alpha$ and $\beta$ are Lagrange multipliers. If one assumes optimal non-singular coding (with $N\gg 1)$, then $l\right>$ becomes $\log_{N}i\right>$ and the solution to the maxent problem is Zipf’s law (Eq. 7) as already pointed out by Mandelbrot [13]. If one assumes $N=1$ or replaces $l\right>$ by the expected rank, that is simply, $i\right>$, it is well-known that the solution will be an exponential distribution [46]. Mandelbrot’s classic argument raises two concerns. The first one is the validity of the maxent argument for non-equilibrium systems such as language. One can argue that the maxent still provides the most unbiased distribution satisfying given constraints [46]. The second one is the validity of the assumption of optimal non-singular coding. Recall that $\eta^{ns}$ is on average $30\%$ [48]. Here we have put forward a class of languages that overcomes the limitations of Mandelbrot’s argument. The members of that class do not need to be coding optimally for Zipf’s law to surface, they only need to be linearly close to the optima.

### v4

@@ 1 Introduction
Table 1: Distinct encodings of numbers from 1 to 6 with an alphabet formed by ${a,b}$.
Code
Unit Non-singular Not non-singular Elias gamma
$1$ $aa$ $aa$ $b$
$2$ $ab$ $aa$ $aba$
$3$ $a$ $a$ $abb$
$4$ $b$ $b$ $aabaa$
$5$ $ba$ $ba$ $aabab$
$6$ $bb$ $bb$ $aabba$
In standard information theory, one assumes that $l(i)$ is the length of the code assigned to unit $i$. The code is a string of characters from some alphabet of size $N$. The problem of compression consists of finding the codes whose length minimizes $l\right>$. We use $l\right>_{min}$ to refer to the minimum value of $l\right>$. If no constraint is imposed on the possible codes, the minima of $l\right>$ are singular. If $l(i)=0$ is allowed, $l\right>$ is minimized by assigning the empty string to each unit. Then $l\right>_{min}=0$ and information transmission is impossible with empty strings. If $l(i)>0$ is required, $l\right>$ is minimized by assigning a string of just one character to each unit. Then $l\right>_{min}=1$ and $N$ has to be as large as the number of units for successful decoding, which, if possible, is eventually an overkill. For this reason, information theory is concerned about constraints on the mapping of units into codes, which are known as coding schemes [11]. In the non-singular coding scheme, every unit is assigned a distinct string. Suppose that $N=2$ and the alphabet is $\{a,b\}$. Table 1 shows a coding table that is non-singular and a coding table that is not non-singular because the string $aa$ is assigned to both unit 1 and unit 2. In the uniquely decodable coding scheme, a particular case of non-singular coding, the segmentation of a string resulting from the concatenation of codes must be unique given the coding table. For instance, $baba$ allows for many interpretations according to the non-singular encoding in Table 1: the segmentation $ba$, $ba$ implies that the units coded are $5$, $5$; the segmentation $b$, $a$, $b$, $a$ implies that the units coded are $4$, $3$, $4$, $3$, and so on. Therefore, the non-singular codes in Table 1 do not satisfy the uniquely decodable coding scheme. In contrast, the Elias gamma encoding [12] in Table 1 does so: the sequence $baba$ can only be segmented as $b$, $aba$ and then the units coded are certainly $1$, $2$.
@@ 2 A new class of coding optimality
The following property indicates that, when Zipf’s law holds, the optimal code length under unique decodability is shifted linearly from the optimal code length under non-singular coding. By the same token, $H$ is shifted linearly from $\log_{N}i\right>$. The same applies to the coding efficiencies.
@@
Zipf’s rank-frequency law with exponent $\alpha$ and normalization factor $c$ is equivalent to a linear dependence between optimal code lengths with parameters $\alpha$ and $\beta=-\log_{N}c$ and $\alpha,\beta\geq 0$. First, $l_{min}^{ud}(i)$ and ł$\l_{min}^{ns}(i)$ are linearly dependent as
$l_{min}^{ud}(i)=\alpha\l_{min}^{ns}(i)+\beta.$
@@
(8, Eqs. 9 and $\sum_{i}p(i)=1$)
Dividing by $l\right>$ on the last result, one obtains Eq. 11. We have $\alpha\geq 0$ by definition of rank. To conclude that $\beta\geq 0$, one does not need to know the precise definition of $c$. Notice that $c=p(1)$ (Eq. 7). As $0<p(1)\leq 1$, $\beta>0$ follows. ∎
@@ 5.1 A new interpretation of rank distributions in log-log scale
Our information theoretic framework provides a new insight into plots of frequency versus rank in double logarithmic scale: the appearance of a straight line indicates (linear) proximity between the lengths of two kinds of optimal codes for a word type given its frequency and its rank (Property 1). The two kinds of optimal codes are the non-singular codes and a subclass that are the uniquely decodable codes. The exponent of Zipf’s law is the slope of the linear transformation that links the optimal lengths. Such an insight does not make any assumption: it just follows from the mathematical shape of Zipf’s law and the optimal lengths of the codes. In a linguistically diverse sample of languages, $\alpha$ ranges from 0.765 in Thai to 1.442 in Korean and has an average value of 1.130 [17]. When controlling for linguistic family, $\bar{\alpha}$, the average $\alpha$ within a linguistic family, exhibits a similar range of variation, i.e. $\bar{\alpha}\in[0.765,1.357]$.
A more powerful insight can be obtained by assuming that the system is compressing, namely reducing the length of codes. In that case, the straight line in log-log scale would indicate proximity of the system to optimal coding. If the gap between the length of words and the corresponding optimal length is only linear, then the system must exhibit Zipf’s law (Property 2). Languages (at least Catalan, Spanish and English) exhibit such a linear proximity, which is consistent with a strong form of compression. If one assumes that all languages belong to that class, one finds that $\bar{\alpha}>1$ in 23 out of 28 families [17], suggesting that $a_{ns}>a_{ud}$ in most families and that $a_{ns}\approx 1.13a_{ud}$ on average across languages (recall Eq. 16). The actual values of $a_{ns}$ and $a_{ud}$ in a diverse sample of languages should be the subject of future research. The failure to find Zipf’s law in certain species but an exponential distribution instead [24, 26, 25] suggests that pressure for compression may not be strong enough for these species or that we are not looking at the right units. In humans, part-of-speech tags, word orders, colors, kinship terms, and verbal alternation classes also exhibit an exponential distribution [35, 36, 28, 37].
@@ 5.3 The origins of Zipf’s law
which implies that rank distributions are channeled to show a power-law (Eq. 7) with $\alpha=1$. However, we have reviewed above multiple real examples of systems that exhibit an exponential rank distribution. Another question is if random typing could be the reason why human languages and other systems show Zipf’s law. In the past, random typing has been regarded as a proof that Zipf’s law can be retrieved without involving any optimization [21, 22, 43, 44]. That argument is flawed because random typing is an optimal non-singular coding system. However, one can still retain that random typing is a proof that Zipf’s law is easy to retrieve by virtue of some simple stochastic process. However, this raises the question of why certain species do not exhibit Zipf’s law for word frequencies. The only answer that random typing provides is that $N=1$ yields an exponential distribution while $N>1$ yields a rank distribution and a frequency spectrum whose ability to mimic the actual distribution of word frequencies has been criticized heavily [45, 46]. Random typing with equal letter probabilities fails to cover the actual range of variation of $\alpha$, e.g., [17], because it only reproduces $\alpha\geq 1$ [47].  [^3] It is well-known that $\alpha=-\log_{N}(1-p_{s})+1\geq 1$ [20]. Notice that such a result can also be retrieved from $\alpha=a_{ns}/a_{ud}=-1/a_{RT}$. A discussion of the potential range of variation of the exponent covered by random typing with unequal character probabilities [20, 48] is beyond the scope of this letter.
@@
where $\alpha$ and $\beta$ are Lagrange multipliers. If one assumes optimal non-singular coding (with $N\gg 1)$, then $l\right>$ becomes $\log_{N}i\right>$ and the solution to the maxent problem is Zipf’s law (Eq. 7) as already pointed out by Mandelbrot [13]. If one assumes $N=1$ or replaces $l\right>$ by the expected rank, that is simply, $i\right>$, it is well-known that the solution will be an exponential distribution [50]. Mandelbrot’s classic argument raises two concerns. The first one is the validity of the maxent argument for non-equilibrium systems such as language. One can argue that the maxent still provides the most unbiased distribution satisfying given constraints [50]. The second one is the validity of the assumption of optimal non-singular coding. Recall that $\eta^{ns}$ is on average $30\%$ [41]. Here we have put forward a class of languages that overcomes the limitations of Mandelbrot’s argument. The members of that class do not need to be coding optimally for Zipf’s law to surface, they only need to be linearly close to the optima. Modern optimization models about the origins of Zipf’s law have similar limitations [47, 38]. For instance, the proposal of a general relevance-resolution trade-off [38] (see Appendix A of [51] for a language-oriented summary of the framework) suggests, at first glance, an alternative general explanation to the origins of Zipf’s law that does not involve word lengths as ours. However, it assumes optimal uniquely decodable encoding when defining speaker and hearer costs and in particular when defining the cost of a word. Recall that $\eta^{ns}$ is on average $40\%$ [41]. In contrast, our proposal does not impose neither optimal coding nor a specific coding scheme. Instead, it uses optimal non-singular coding and optimal uniquely decodable encoding as reference points from which natural languages are hypothesized to deviate linearly from. While we aim to provide a unified explanation to linguistic laws via compression in communication systems, [38] has a much wider scope in terms of systems. We hope that our research stimulates a dialogue between these two complementary approaches.